1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

The Werewolf Varulven La lupfanto- mo The Banshee (An Approach)

tradukita de Alexander Gross tradukita de Sven Collberg tradukita de Bertilo Wennergren tradukita de Max Knight

A Werewolf, troubled by

his name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

En varulv lopp från

fru m.m. en vacker afton samt

begav sig till en bymagi-

sters grav och bad: “Var snäll

och konjugera.

Jen lupfantom’ de famili’ meznokte iam ku-

ris for al tombo de la pro-

fesor’. kaj petis pri konjugaci’.

One night, a banshee slunk away

from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect

me, pray.”

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready. mig!” Byskolläraren

stod upp på bleckna-

mnsskyltens

mässingknopp med svar till den,

som satt på pass med tåligt korslagd

tass på tass.

Do diktis tiu dok-

ta ul’ starante sur la kructabul’ dum lup’ senmova kiel splinto atentis vortojn de l’ mortinto.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition

said who meekly knelt before the dead:

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

“Jag var-ulv 1:a pers.

i sing., du, han, hon, den,

det var-ulv. Så vi voro-ulv, I voren.

På de voro-ulv kom in-

genting.”

De “lupfantomo”, simpla rekta, tra “lupfintomo”, plej perfekta, plu “lupfontomo”, iom rara, ĝis “lupfuntomo”, ho, koŝmara.

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that

is all.”

“I know that–I’m no mental

cripple– The future form and parti-

ciple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

Varulven myste glad

åt hur pers. hade följts av

sing. och plur. “Fast”, bad han, “kan

det inte ges ens en enda pers. i tem-

pus presens?”

Ĉi participoj plaĉis multe al la fantomo, kaj rezulte ĝi petis: “Legu plu el PIV-o pri participoj en pasivo.”

The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not

add, so far, the plural to the singu-

lar!”

…  …  …  …

2/3

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-

wolf?’ Utterly wrong! Such words are wounds

beyond all suture– I’m sorry, but you have no

future.”

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine im-

perfecto var contradictio in

adjecto.

La viro ĉagreniĝis tamen: “Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

The teacher, though, admitted then that this was not within

his ken. “While bans are frequent”, he advised, “A she cannot be pluri-

zed.”

The Werewolf knew better–

his sons still slept At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste exi-

stera.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

The banshee, rising

clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu  la  re-

tejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .